OpenStudy (anonymous):

Find the measure of each interior angle of a regular octagon. (1 point) (0 pt) 135° (0 pts) 1,080° (0 pts) 180° (0 pts) 100°

OpenStudy (anonymous):

iknow its 132 i just need to know how can you get the answer in an example

OpenStudy (anonymous):

135

OpenStudy (mathstudent55):

The sum of the interior angles of an n-sided polygon is (n - 2)180. In a regular polygon, all angles interior are congruent. Use the formula above with n = 8 Then divide answer by 8

OpenStudy (anonymous):

can you show step by step to get the answer please?

OpenStudy (mathstudent55):

(n - 2) 180 An octagon has 8 sides, so n = 8 (8 - 2)180 = 6(180) = 1080 1080 is the measures of all angles added up together. Since there are 8 congruent (same measure) angles, each angle measures 1080/8 = 135